the resistance of copper wire varies directly with its length. a copper wire 275 feet long has a resistance of 950 ohms. what is the resistance of 1575 feet of copper wire
Answer by stanbon(48564)

the resistance of copper wire varies directly with its length.
R = k*L
--------------
a copper wire 275 feet long has a resistance of 950 ohms.
Solve for "k":
950 = k*275
k = 975/275 = 38/11
-----------------------
Equation: R = (38/11)*L
------------
what is the resistance of 1575 feet of copper wire
R = (38/11)*1575
Resistance = 5440.9090... ohms
==================================
Cheers,
Stan H.
---------------------------